Consider a monopolistically competitive market with N firms. Each firm's business opportunities are described by the following equations:


Demand: Q=100/N−P


Marginal Revenue: MR=100/N−2Q


Total Cost: TC=50=Q2


Marginal Cost: MC=2Q


a. How does N, the number of firms in the market, affect each firm's demand curve? Why?


b. How many units does each firm produce? (The answers to this and the next two questions depend on N.)


c. What price does each firm charge?


d. How much profit does each firm make?


e. In the long run, how many firms will exist in this market?

1
Expert's answer
2022-03-25T15:06:56-0400

a)The more number of firms, the flatter their demand curves. Because the demand equation shows the inverse relationship between Q and N.

b) Each firm produces at the point where MR=MC, so 100/N-2Q=2Q; 100/N=4Q; Q=100/4N

c) The price can be found from the demand equation :100/N-P=Q=100/4N; P=100/N-100/4N=(400-100)/4N=300/4N

d) Profit is PQ-TC=100/4N×300/4N-50=1875/N^2 -50

e) In the long run the quantity of firms will increase, the price will decrease, and the firms will produce at the break-even point